OpenStudy (anonymous):

Simplify √36x^7 How would you go about this? Please show the work and explain it so I can understand. Thank you!

OpenStudy (anonymous):

\[\sqrt{a \times b} = \sqrt{a} \times \sqrt{b}\]

OpenStudy (anonymous):

\[\sqrt{x} = x^{1/2}\]

OpenStudy (anonymous):

\[(x^a)^b = x^{a \times b}\]

OpenStudy (anonymous):

Ah ok i think ive got it, thanks.